OpenStudy (accidentalaichan):

1/5 (x + 2)= 8 Will give medal! ^^

OpenStudy (anonymous):

Solve for x: (x+2)/5 = 8 Multiply both sides by a constant to simplify the equation. Multiply both sides of (x+2)/5 = 8 by 5: (5 (x+2))/5 = 5×8 Cancel common terms in the numerator and denominator of (5 (x+2))/5. (5 (x+2))/5 = 5/5×(x+2) = x+2: x+2 = 5×8 Multiply 5 and 8 together. 5×8 = 40: x+2 = 40 Isolate terms with x to the left hand side. Subtract 2 from both sides: x+(2-2) = 40-2 Look for two terms that sum to zero. 2-2 = 0: x = 40-2 Evaluate 40-2. 40-2 = 38: Answer x = 38

OpenStudy (accidentalaichan):

Thanks for breaking it down ^^

OpenStudy (anonymous):

You're welcome.